OpenStudy (anonymous):

Please help!! Part A: Two similar prisms have heights of 4cm and 10cm. What is the ratio of their surface areas? A. 2:5 B. 4:25 C. 8:125 D. 16:625

OpenStudy (joann88):

A

OpenStudy (anonymous):

Could you explain please?

OpenStudy (joann88):

Because their heights have the ratio 2:5 and the prisms are similar. That's why their surface have the same ratio - 2:5.

OpenStudy (anonymous):

Could I ask you another one like it?

OpenStudy (joann88):

Go on

OpenStudy (anonymous):

What is the ratio of the volumes of the prisms in the previous problem? Would it be the same answer?

OpenStudy (anonymous):

The answer choices are the same also

OpenStudy (anonymous):

My book says that the ratio for volumes is a^3:b^3 and the ratio for areas is a^2:b^2. Sorry idk if that helps...

OpenStudy (joann88):

Oh, I'd made mistake.. I didn't mention that prisms are stereometrical. Ok.. Step by step. Prism have square surface area, right? And each side of SurfArea of one prism have the ratio with the same side of the other prism 2:5. So, surface areas have ratio 4:25. Now their volumes must have ratio 8:125, because each of three multipliers in volume_formula have the ratio 2:5. Sorry, for first mistake)

OpenStudy (anonymous):

That's okay. The second one is 8:125 right?

OpenStudy (joann88):

Yes

OpenStudy (anonymous):

I'll give you a medal if you help with some more??

OpenStudy (joann88):

Let's try)

OpenStudy (anonymous):

Which of the following could represent the scale factor of the larger cylinder to the smaller cylinder? Large ones diameter = 18, height = 30. Smaller ones diameter = 12, height = 20.

OpenStudy (anonymous):

Answer choices(pick two) A. 15:9 B. 5:3 C. 3:2 D. 21:14 E. 2:3 F. 20:30

OpenStudy (joann88):

Well... If I'm understanding correctly, if we multiply larger cylinder by scale factor then we must receive small cylinder, right? Sizes of smaller cylinder have the ratio with sizes of larger one 2:3. So, we must take scale factor 2:3 or 20:30 in order to make correct calculation.

OpenStudy (anonymous):

So it's 2:3 and 20:30?

OpenStudy (joann88):

Yes

OpenStudy (anonymous):

A cylinder with a 4 in. Diameter and a 6 inch height holds 1 pound of oatmeal. To the nearest ounce, how much oatmeal will a similar 10 inch tall cylinder hold? (Hint: 1 pound = 16 ounces)

OpenStudy (joann88):

I will answer later. I must run)

OpenStudy (anonymous):

Answer choices: A. 37 B. 44 C. 74 D. 192 I'm guessing it's 37 but not sure

OpenStudy (anonymous):

Could I talk to you tonight at like 8:30?

OpenStudy (onyxxelite):

hello

OpenStudy (joann88):

Now it is 22:36 in the city, where I am. If we have similar time or you have earlier time - it will be Ok. Otherwise, I can't guarantee that I will answer)

OpenStudy (joann88):

I've got 74. If cylinders are similar then their sizes have ratio 3:5 (height1/height2=6/10=3/5). That's why their volumes have ratio (3/5)^3. So, vol1/vol2 = 27/125, and vol2 = 125/27*1pound*16ounces.

OpenStudy (dopeness_):

Pease close this post and post math questions in the math section.